OpenStudy (anonymous):

show that q /(x^2-2) is not isomorphic to q /(x^2-3)

terenzreignz (terenzreignz):

q here being the set of rational numbers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Q[x] /(x^2-2) , Q[x] /(x^2-3), the congruence classes

terenzreignz (terenzreignz):

Also, did you note that \[\huge \frac{\mathbb Q}{<x^2-2>}\cong\mathbb{Q}(\sqrt2)\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

oh and Q(sqr(3) is isomorphic to Q / x^2 - 3

terenzreignz (terenzreignz):

And, also, \[\huge \frac{\mathbb{Q}}{<x^2-3>}\cong\mathbb{Q}(\sqrt3)\]yes.

OpenStudy (anonymous):

so i have all the pieces, i just dont know how to finish the puzzle

terenzreignz (terenzreignz):

Well, here's one way to do it, suppose that there is an isomorphism from Q(sqrt2) to Q(sqrt3)

terenzreignz (terenzreignz):

Let that isomorphism be \[\huge \Phi:\mathbb Q(\sqrt2) \ \rightarrow \ \mathbb Q(\sqrt3)\]

terenzreignz (terenzreignz):

So, what's \[\large \Phi(1)\]?

terenzreignz (terenzreignz):

Unity must be mapped to unity, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

But I do not understand what \[\Phi(1) is\]

terenzreignz (terenzreignz):

It's just the image of 1, under the isomorphism, phi. If you don't like phi, we can use f instead. f(1) = 1 right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

So, by extension, the image of any integer must be itself, since, for any integer n, f(n) = f(1+1+1...+1) = f(1) + f(1) + f(1) ... + f(1) = 1+1+1+...+1 = n

OpenStudy (anonymous):

I follow so far.

terenzreignz (terenzreignz):

So, the image of any integer is itself, what does that make the image of any rational number?

OpenStudy (anonymous):

itself?

terenzreignz (terenzreignz):

Remember, since it's a field isomorphism, \[\huge f\left(\frac{m}n\right)=\frac{f(m)}{f(n)}\]

OpenStudy (anonymous):

but sqrt of 2 and 3 arent rational?

terenzreignz (terenzreignz):

We'll get to those. So, we're clear that the image of any rational number must be itself, okay? :)

OpenStudy (anonymous):

haha yes

terenzreignz (terenzreignz):

Okay, so, what must be \[\huge f(\sqrt2)\]?

terenzreignz (terenzreignz):

Relax, we don't know :) But whatever it is, it must not be rational. Right?

OpenStudy (anonymous):

Yes

terenzreignz (terenzreignz):

so... \[\huge f(\sqrt2)=a+b\sqrt3\]where a and b are rational, and b is not equal to zero.

OpenStudy (anonymous):

Because Qsqrt 2 is iso to Qsqrt3

OpenStudy (anonymous):

based on our supposition

terenzreignz (terenzreignz):

Not that, yet, but only the fact that f maps Q(sqrt2) to Q(sqrt3)

OpenStudy (anonymous):

oh okay

terenzreignz (terenzreignz):

So, based on this, what now is \[\huge f(\sqrt2^2)=f(2)\]?

terenzreignz (terenzreignz):

We know that since 2 is rational, its image must also be 2, right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

And since this is an isomorphism, it is also equal to \[\huge (a+b\sqrt3)^2=a^2+2ab\sqrt3+3b^2\]

OpenStudy (anonymous):

Yup

terenzreignz (terenzreignz):

So, \[\huge 2=a^2+2ab\sqrt3+3b^2 \]see anything wrong yet?

terenzreignz (terenzreignz):

The left side is obviously rational, so it forces the right side to also be rational. So, 2ab must be 0, or ab must be zero.

OpenStudy (anonymous):

but didnt we say b wasn't zero?

terenzreignz (terenzreignz):

That's right, so a must be zero :)

terenzreignz (terenzreignz):

And so, \[\huge 2=3b^2\]

OpenStudy (anonymous):

b = sqrt of 2/3

terenzreignz (terenzreignz):

which is not rational. Contradiction, anyone? :)

OpenStudy (anonymous):

haha yes! because f(2) forces the other side to be rational?

terenzreignz (terenzreignz):

Yes. So, satisfied? ^.^

terenzreignz (terenzreignz):

Because assuming that Q(sqrt2) is isomorphic to Q(sqrt3) leads us to sqrt(2/3) is rational, a contradiction.

terenzreignz (terenzreignz):

Make your presence felt, @urbanderivative :)

OpenStudy (anonymous):

I was just reading

OpenStudy (anonymous):

I get it now, thanks for all your help!

terenzreignz (terenzreignz):

Okay, that's all I need to know :) I need to go now, thanks for the nice question ^.^ ---------------------------------------------- Terence out

OpenStudy (anonymous):

I like to keep you sharp